OpenStudy (alyssajobug):

Algebra 2 help?!?!?! The height h in feet of a baseball on Earth after t seconds can be modeled by the function h(t) = -16(t - 1.5)2 + 36, where -16 is a constant in ft/s2 due to Earth's gravity. The gravity on Mars is only 0.38 times that on Earth. If the same baseball were thrown on Mars, it would reach a maximum height 59 feet higher and 2.5 seconds later than on Earth. Write a height function for the baseball thrown on Mars. Help? Now sure how to adjust the equasion...

OpenStudy (alyssajobug):

Would it be just making it -16t^2+24+30=0 turned into y = -6.08(t - 4)2 + 95 by multiplying by 0.38?

OpenStudy (phi):

looks good

OpenStudy (phi):

I would do the problem by noticing the vertex on earth is (1.5,36) the vertex on mars will be (1.5+2.5, 36+59) or (4,95) the -16 coefficient on earth gets changed to -16*0.38= -6.08 so in vertex form, plugging in the numbers for (h,k) and a in y= a(x-h)^2 + k y= -6.08(t-4)^2 + 95

OpenStudy (alyssajobug):

ok, great so I did it right? so will my way of just multiplying each thing by 0.38 work all the time or was this just a fluke?

OpenStudy (phi):

good question.

OpenStudy (phi):

the earth's equation comes from physics, and starts with earth's force due to gravity... so your trick probably works for this situation. But if the equation were some random parabola, I don't think it works. I think they expect you to solve it the way I did.

OpenStudy (alyssajobug):

Oh, ok I'll copy and paste your comment to a word where I can use it as an example then! THANKS!! AGAIN!!!

OpenStudy (phi):

can you explain what you did in more detail ?

OpenStudy (alyssajobug):

I took the original equasion (h(t) = -16(t - 1.5)^2 + 36) and multiplied each section by 0.38 (h(t) = -16x0.38 (t - 1.5x0.38)^2 + 36 to get y = -6.08(t - 4)^2 + 95 as the only option that had both of the first two right, not sure how I got from 36 to 95, my notes are beyond sloppy but I did it some how...

OpenStudy (phi):

ok, in that case, the only thing you should multiply by 0.38 is the -16. you "tweak" the 1.5 and the 36 by adding the numbers they told you.

OpenStudy (alyssajobug):

oh ok so 2.5 s longer and 59 ft higher gets added there?

OpenStudy (phi):

yes